Remy walked to a friend's house, m miles away, at an average rate of 4 mph. The m-mile walk home was only at 3 mph. Remy spent 2 hours walking in all. Find the value of m.
Answer by stanbon(75887)

Remy walked to a friend's house, m miles away, at an average rate of 4 mph. The m-mile walk home was only at 3 mph. Remy spent 2 hours walking in all. Find the value of m.
-----
Going DATA: distance = m miles ; rate = 4 mph ; time = d/r = m/4 hours
----
Return DATA: distance = m miles ; rate = 3 mph ; time = d/r = m/3 hours
------
Equation:
time + time = 2 hrs
m/4 + m/3 = 2
3m + 4m = 24
7m = 24
m = 24/7 = 3 3/7 miles
=======================
Cheers,
Stan H.